OpenStudy (anonymous):

factor completely: 360v^3-702v^2+324v=18v

OpenStudy (anonymous):

some problem here because you wrote an equation with \[=18v\] on the right hand side. is that a typo?

OpenStudy (anonymous):

no no typo

OpenStudy (anonymous):

move the 18 v over and set it equal to 0 solve for a GCF and then factor the trinomial square

OpenStudy (anonymous):

360v^3 -703v^2 +324v -18v =18v-18v=0

OpenStudy (anonymous):

so what am i left with ?

OpenStudy (anonymous):

360v^3-703v^2+306v=0 coefficients look coprime might have to synthetic divide are you familiar with this technique?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh my mistake its 702 v^2 then take out an even greatest common factor

hero (hero):

I thought you were done with this factoring stuff

OpenStudy (anonymous):

not to be annoying but if it is set equal to something then the question should be "solve for v" not "factor"

OpenStudy (anonymous):

v will have an actual value (or a choice of values) here. maybe you can solve by factoring

OpenStudy (anonymous):

@hero i thought so too

OpenStudy (anonymous):

lets factor out 18v

OpenStudy (anonymous):

now i am wondering if it really did mean "factor" and ignore the right hand side, because the left hand side factors nicely as \[18 v (4 v-3) (5 v-6) = 18 v\]

OpenStudy (anonymous):

whereas if you subtract 18v from both sides you get nothing very nice. where did the question come from?

OpenStudy (anonymous):

a quiz

OpenStudy (anonymous):

and it just said "factor" not "solve for v"?

OpenStudy (anonymous):

yea it says factor completley

OpenStudy (anonymous):

well tell your math teacher that if you have an equation, the only way for it to be true is if v takes on certain values. it really doesn't make sense to give an equation (something with an equal sign, like \[x^2-2x+1=5\] and say "factor"

OpenStudy (anonymous):

like i guess paraenthisese are supposed to go after 18v

OpenStudy (anonymous):

we know the obvious answer is v=0 the other factors are complex you could use the quadratic formula but i am too lazy :)

OpenStudy (anonymous):

i need the quaratic formula

OpenStudy (anonymous):

when given an quadratic equation of the form ax^2 +bx+c=0 where a,b,c are any elements of the reals, you can use x=(-b +/- sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

nevermiond lol i go it

hero (hero):

You do?

OpenStudy (anonymous):

what about 15h^2-17hj-4j^2

hero (hero):

I'm done with doing these for right now

hero (hero):

Oh, and Maria. Just a word of advice, you should get rid of the middle finger.

OpenStudy (anonymous):

how about tht

hero (hero):

about what?

OpenStudy (anonymous):

the new picture

hero (hero):

It's still the same from my end.

hero (hero):

Oh nice. Much better :o)